OpenStudy (anonymous):

x = 7 * a a = 9 * 7 / 2 x * a

OpenStudy (anonymous):

Please help

OpenStudy (anonymous):

You well get medal if you help

OpenStudy (anonymous):

SOMEONE JUST HELP PLEASE

OpenStudy (anonymous):



jhonyy9 (jhonyy9):

x=7*a a = 9*7/2 x*a = (7a)*a = (7*9*7/2) *(9*7/2) = (7*9*7*9*7)/(2*2) = ?